You have calculated via OLS: y = -1.29 + x*0.0452 Estimate the marginal effect as x goes from 1 to 2
Answer #1

Given,

y = -1.29 + x*0.0452

Putting, x = 1

y = -1.29 + (1)*0.0452

= -1.2448

So, At x = 1, y = -1.2448

Now, putting x = 2.

y = -1.29 + (2)*0.0452

y = -1.1996

And, At x = 2, y = -1.1996

Marginal effect as x moves from 1 to 2 = (Change in y / change in x)

= (-1.1996 + 1.2448) / (2-1) = 0.0452

Therefore, Marginal effect as x goes from 1 to 2 = 0.0452

Please don't forget to rate the answer if its helpful, thank you.